OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = -8, an = 5 • an-1 a. -8, -40, -200, -1000, -5000, -25,000 b. -8, -40, -35, -30, -25, -20 c. 0, 5, -40, -35, -30, -25 d. -40, -200, -1000, -5000, -25,000, -125,000

OpenStudy (anonymous):

a classic recurcive sequence, answer is a

OpenStudy (anonymous):

thx! how exactly do you know? i just learned this.

OpenStudy (anonymous):

\[a _{1}=-8 \]\[a _{n}=5*a _{n-1}\] assume n=2, for the second term in the sequence, \[a _{2}=5a _{2-1}=5a_1=5*-8=-40\] and so on. So the third term becomes -200; at this point only one answer matches the first three terms.

OpenStudy (anonymous):

okay i get it :) thank you!